OpenStudy (anonymous):

solve log(X^2) - log(x+4) =log12

OpenStudy (anonymous):

You could do log(x^2/x+4)=log(12) and then apply a base 10 exponential to both sides

OpenStudy (anonymous):

Which would cancel out the logs and leave you with x^2/x+4=12

OpenStudy (anonymous):

It says to find exact solutions, so would I then multiply by x+4 and eventually complete the square to find what x would equal?

OpenStudy (anonymous):

Yep